### v4

# Enumeration of Elliptic Curves via Elliptic Gromov-Witten Invariants of Four Dimensional Projective Fano Hypersurfaces

Abstract
In this paper, we propose a conjecture that clarifies the relationship between the number of degree $d$ elliptic curves in complex four-dimensional projective Fano hypersurfaces and their degree $d$ elliptic Gromov-Witten (GW) invariants. The elliptic GW invariants are computed using the elliptic virtual structure constants proposed in our previous works [^1] Keywords; enumerative geometry; gromov-witten invariants; mirror symmetry; elliptic curve; projective Fano hypersurface. 2020 MATHEMATICS SUBJECT CLASSIFICATION; 14N10, 14N35. .
1 Introduction
The genus $g$ and degree $d$ Gromov-Witten (GW) invariants of a complex $n$-dimensional Kähler manifold $X$ are generally expected to count the number of holomorphic curves of genus $g$ and degree $d$ in $X$ that satisfy the passing-through conditions imposed by the operator insertions. Therefore, these invariants are anticipated to be non-negative integers. However, there are many cases where they fail to be non-negative integers. As far as we know, the genus 0 GW invariants of projective Fano manifolds are always non-negative integers. In the case of Calabi-Yau and general-type manifolds, however, genus 0 GW invariants typically become rational numbers due to contributions from multiple cover maps. For Calabi-Yau manifolds, many studies analyze the contributions from multiple cover maps, often relating them to open Calabi-Yau manifolds obtained as vector bundles over $CP^{1}$ [1, 12]. In contrast, for general-type manifolds, there are few works on contributions from multiple cover maps to genus 0 GW invariants, with the exception of the preprint [8] in 2004.
Now, let us turn to the case of genus $1$ GW invariants. The genus $1$ GW invariants of projective Fano surfaces (complex two-dimensional manifolds) are non-negative integers, at least for lower degrees. However, in the Calabi-Yau case, we know that the elliptic GW invariants include not only contributions from multiple cover maps from a genus $1$ curve but also degenerate contributions associated with genus 0 GW invariants [2, 15]. Roughly speaking, the degenerate contributions arise from the boundary loci of the moduli space of stable maps from genus $1$ stable curves, where the stable curves degenerate into nodal curves containing a genus 0 component. At these loci, the stable curves can be mapped into rational curves in the target manifold $X$, and the contributions are related to the genus 0 GW invariants of $X$. It is well-known that degenerate contributions also appear in the case of complex three-dimensional Fano manifolds. Heuristically, this phenomenon is expected from the dimensional counting of the virtual dimension of the moduli space of stable maps. The virtual dimension of the moduli space of stable maps from stable curves to complex three-dimensional projective manifolds with fixed degree $d$ is independent of the genus of the stable curves. Hence, we can speculate that the genus $1$ GW invariants may contain degenerate contributions because the operator insertions required to balance the positive virtual dimension of the moduli space do not depend on the genus. A natural question thus arises: **”Do degenerate contributions appear in the case of a projective Fano manifold whose complex dimension is greater than 3?”**
In this paper, motivated by the above question, we investigate the relationship between the elliptic GW invariants of four-dimensional projective Fano manifolds and the number of elliptic curves. In the case of complex four-dimensional Kähler manifolds, the virtual dimension of the moduli space of stable maps from genus 0 stable curves is one complex dimension greater than that from genus $1$ stable curves. Therefore, it is not clear whether degeneration contributions appear, based on the naive speculation used in the preceding discussion. In the Calabi-Yau case, the work [14] by Klemm and Pandharipande shows that degenerate contributions do indeed appear, and they are related to genus 0 GW invariants in a more complicated way than in the case of complex three-dimensional Calabi-Yau manifolds. We must note that their results rely on the computation of genus $1$ GW invariants using the BCOV conjecture [2]. On the other hand, in [7], we proposed a conjecture that enables a kind of B-model computation (from the perspective of mirror symmetry) of genus $1$ GW invariants for projective hypersurfaces of arbitrary dimension. We therefore used our conjecture to evaluate genus $1$ GW invariants of four-dimensional projective Fano hypersurfaces. In the Fano case, the Virasoro conjecture [4] and Getzler’s equation [6] are also effective for this purpose. When applying these methods to evaluate genus $1$ invariants, one must use information from genus 0 GW invariants with insertions of operators derived from primitive cohomology classes of the hypersurfaces. As was done in [5], these genus 0 invariants were computed in the case of a cubic hypersurface in $CP^{4}$ using the associativity equation, since it has primitive classes only in the bidegree $(2,1)$ and $(1,2)$ sectors. However, in the four-dimensional case, a projective Fano hypersurface can have primitive classes in the bidegree $(3,1)$, $(2,2)$, and $(1,3)$ sectors. Therefore, we must overcome technical obstacles to compute genus 0 invariants with these operator insertions using known techniques. Moreover, as we experienced in the computation of genus $1$ invariants of the cubic hypersurface using the Virasoro conjecture [5], the method requires outrageously complicated and lengthy computations. In contrast, our conjecture does not require any information regarding genus 0 GW invariants with insertions of primitive classes and can be easily automated using computer software such as Maple and Mathematica.
After computing the genus $1$ GW invariants of the degree $k$ ($k=1,2,3,4,5$) hypersurface in $CP^{5}$ (denoted by $M_{6}^{k}$) up to degree $d=5$, we concluded that degeneration contributions do appear because the invariants often turn out to be negative rational numbers. We then investigated how these contributions are described by the genus 0 GW invariants of the hypersurface. The guiding principles we used for this determination were:
  1. 1.
    The number of elliptic curves of degree $d=1,2$ in $M_{6}^{k}$ that satisfy the passing-through conditions imposed by operator insertions is always 0.
  2. 2.
    The number of elliptic curves of degree $d=3$ in $M_{6}^{k}$ that satisfy the passing-through conditions imposed by operator insertions can be computed by applying the method presented in the preprint [11] by S. Katz.
  3. 3.
    The expected formula is considered an extension of the result presented in [14] to the Fano case.
Using these hints, we constructed a conjectural formula that describes the degenerate contributions in terms of genus 0 GW invariants, which will be presented in the next section. Heuristically, the genus $1$ GW invariants of degree $d$ are considered the sum of the degenerate contributions and the number of elliptic curves of the same degree satisfying the passing-through conditions. Therefore, if we subtract the degenerate contributions from the genus $1$ GW invariants, we expect to obtain non-negative integers. We denote the resulting number by $E_{d,a,b,c}$, and it turns out to be always a non-negative integer up to $d=5$. Specifically, $E_{d,a,b,c}=0$ if $d=1,2$, and it coincides with the number of elliptic curves computed by S. Katz’s method if $d=3$. The list of $E_{d,a,b,c}$ for $M_{6}^{k}$ ($k=1,2,3,4,5$) up to $d=5$ will be presented in Section 3. With these results, we propose a conjectural method for counting elliptic curves in a four-dimensional projective Fano hypersurface via genus $1$ GW invariants.
2 Definitions and Our Main Conjecture
Let $M_{N}^{k}$ denote a degree $k$ hypersurface in $CP^{N-1}$. Let $h$ be the pull-back of the hyperplane class of $H^{1,1}(CP^{N-1},{\bf C})$ to $H^{1,1}(M_{N}^{k},{\bf C})$ via the inclusion map $\iota:M_{N}^{k}\hookrightarrow CP^{N-1}$. We denote the genus $g$ and degree $d$ Gromov-Witten (GW) invariants of $M_{N}^{k}$ by $\langle\prod_{j=1}^{L}({\cal O}_{\alpha_{j}})^{m_{j}}\rangle_{g,d}$, where $\alpha_{j}$’s ($j=1,\cdots,L$) are a linear basis of $H^{*,*}(M_{N}^{k},{\bf C})$. We omit the rigorous definition of the Gromov-Witten invariant here. In this paper, we only consider the following GW invariants:
$\langle\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{0,d}\quad(g=0,1).$ (2.1)
In the $g=0$ and $g=1$ cases, it is well-known that the GW invariants are non-zero only if the following conditions are satisfied:
$\langle\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{0,d}\neq 0\;\Longrightarrow\sum_{a=0}^{N-2}m_{a}(a-1)=N-5+d(N-k),\;\;$ (2.2)
$\langle\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{1,d}\neq 0\;\Longrightarrow\sum_{a=0}^{N-2}m_{a}(a-1)=d(N-k).\;\;$ (2.3)
For later use, we also introduce the **Kähler (divisor) equation** and the **puncture equation** for the GW invariants.
$\langle{\cal O}_{h}\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}=d\langle\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}\quad(d\geq 1),$ (2.4)
$\langle{\cal O}_{h^{0}}\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}=0\hskip 105.2751pt(d\geq 1).$ (2.5)
If $d=0$, the genus 0 GW invariants are non-vanishing only if the number of operator insertions $\sum_{a=0}^{N-2}m_{a}$ equals $3$, and the non-vanishing invariants are given by:
$\langle{\cal O}_{h^{a}}{\cal O}_{h^{b}}{\cal O}_{h^{c}}\rangle_{0,0}=\int_{M_{N}^{k}}h^{a}\wedge h^{b}\wedge h^{c}=k\cdot\delta_{a+b+c,N-2}.$ (2.6)
If $g=1$ and $d=0$, the only non-vanishing GW invariant is $\langle{\cal O}_{h}\rangle_{1,0}$, which is given by
$\langle{\cal O}_{h}\rangle_{1,0}=-\frac{1}{24}\int_{M_{N}^{k}}h\wedge c_{N-3}(T^{\prime}M_{N}^{k}),$ (2.7)
where $T^{\prime}M_{N}^{k}$ is the holomorphic tangent bundle of $M_{N}^{k}$ and $c_{N-3}(T^{\prime}M_{N}^{k})$ is its second-to-top Chern class.
From now on, we fix $N=6$. Hence, we only consider $M_{6}^{k}$, the four-dimensional hypersurface in $CP^{5}$. Considering the Kähler equation and the puncture equation, we can easily see that the non-trivial $g=0,1$ GW invariants with positive degrees are given by
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{0,d},\quad\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{1,d}\quad(d\geq 1)$ (2.8)
From (2.3), all $g=1$ and $d\geq 1$ GW invariants (given by (2.8)) vanish if $k>6$. In the $k=6$ case, the non-trivial genus 0 and $1$ GW invariants are given by $\langle{\cal O}_{h^{2}}\rangle_{0,d}$ and $\langle*\rangle_{1,d}$ ($d\geq 1$), respectively (* means no operator insertions). The relation between $\langle*\rangle_{1,d}$ and the number of elliptic curves of degree $d$ in $M_{6}^{6}$ was extensively studied in [14]. Therefore, we focus on the $k=1,2,3,4,5$ cases. In these cases, we evaluated the genus 0 GW invariants in (2.8) using both the multi-point virtual constants [10] and the associativity equation [13]. We then evaluated the corresponding genus $1$ invariants using the method of elliptic virtual structure constants [7]. Based on these numerical results and the three criteria introduced in Section 1, we propose the following conjecture:
Conjecture 1.
Let $E_{d,a,b,c}$ be defined by the following relation:
$\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{1,d}=-\frac{(k^{2}-6k+15)d-(6-k)}{24d}\langle({\cal O}_{h^{2}})^{a+1}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{0,d}$
$+\frac{1}{2}\sum_{\begin{subarray}{c}0<d_{1},d_{2}\\ d_{1}+d_{2}=d\end{subarray}}\frac{2d-d_{1}d_{2}(6-k)}{24d}\sum_{l=0}^{a}\sum_{m=0}^{b}\sum_{n=0}^{c}\binom{a}{l}\binom{b}{m}\binom{c}{n}$
$\times\left(\frac{1}{k}\langle({\cal O}_{h^{2}})^{l}({\cal O}_{h^{3}})^{m}({\cal O}_{h^{4}})^{n}{\cal O}_{h^{3}}\rangle_{0,d_{1}}\langle{\cal O}_{h}({\cal O}_{h^{2}})^{a-l}({\cal O}_{h^{3}})^{b-m}({\cal O}_{h^{4}})^{c-n}\rangle_{0,d_{2}}\right$.
$\left.+\frac{1}{k}\langle({\cal O}_{h^{2}})^{l}({\cal O}_{h^{3}})^{m}({\cal O}_{h^{4}})^{n}{\cal O}_{h}\rangle_{0,d_{1}}\langle{\cal O}_{h^{3}}({\cal O}_{h^{2}})^{a-l}({\cal O}_{h^{3}})^{b-m}({\cal O}_{h^{4}})^{c-n}\rangle_{0,d_{2}}\right$.
$\left.+\frac{1}{k}\langle({\cal O}_{h^{2}})^{l}({\cal O}_{h^{3}})^{m}({\cal O}_{h^{4}})^{n}{\cal O}_{h^{2}}\rangle_{0,d_{1}}\langle{\cal O}_{h^{2}}({\cal O}_{h^{2}})^{a-l}({\cal O}_{h^{3}})^{b-m}({\cal O}_{h^{4}})^{c-n}\rangle_{0,d_{2}}\right)$
$+\frac{k-2}{48}a!b!c!\langle({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\rangle_{0,1}\delta_{k,5}\delta_{d,2}+E_{d,a,b,c}.$ (2.9)
Here, $\binom{a}{l}$ denotes a binomial coefficient, and $\delta_{i,j}$ denotes the Kronecker delta. Then, $E_{d,a,b,c}$ is expected to be an integer representing the number of elliptic curves.
Remark 1.
The first and second Chern classes of $T^{\prime}M_{6}^{k}$ are given by
$c_{1}(T^{\prime}M_{6}^{k})=(6-k)h,\;\;c_{2}(T^{\prime}M_{6}^{k})=(k^{2}-6k+15)h^{2}.$ (2.10)
Note the appearance of the polynomials in $k$, (6k)(6-k) and $(k^{2}-6k+15)$, in (2.9).
Heuristically, we expect that $E_{d,a,b,c}$ is the number of elliptic curves of degree $d$ in $M^{k}_{6}$ passing through $a$ 33-planes, $b$ 22-planes, and $c$ 11-planes in $CP^{5}**$. In particular, for $d=1,2$, our numerical computation confirmed that $E_{d,a,b,c}$ always vanishes. Furthermore, for $d=3$, we confirmed that $E_{d,a,b,c}$ coincides with the number of elliptic curves satisfying the required passing-through conditions, as computed by applying the method presented in [11] by S. Katz. We will explain the outline of S. Katz’s method in Appendix A.
3 Lists of $g=0,1$ Gromov-Witten Invariants and $E_{d,a,b,c}$ for $M^{k}_{6}$ ($k=1,2,3,4,5$) up to $d=5$
In this section, we present the numerical results used to determine the formula (2.9). For simplicity, we denote the genus $g$, degree $d$ Gromov-Witten (GW) invariants of $M_{6}^{k}$, $\left<({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\right>_{g,d}$, by $N^{g}_{d,a,b,c}$:
$N^{g}_{d,a,b,c}:=\left<({\cal O}_{h^{2}})^{a}({\cal O}_{h^{3}})^{b}({\cal O}_{h^{4}})^{c}\right>_{g,d}.$
The genus 0 GW invariants were computed using the **associativity equation** and the **multi-point virtual structure constants**, while the genus $1$ invariants were computed following the method described in [7].
Table 1: $g=0,M_{6}^{1},d=1,2$
d (a,b,c) $N^{0}_{d,a,b}$
1 (0,0,2) 11
1 (1,1,1) 11
1 (3,0,1) 11
1 (0,3,0) 11
1 (2,2,0) 22
1 (4,1,0) 33
1 (6,0,0) 55
2 (0,1,3) 0
2 (2,0,3) 11
2 (1,2,2) 11
2 (3,1,2) 44
2 (5,0,2) $11$
2 (0,4,1) 22
2 (2,3,1) 66
2 (4,2,1) 2121
2 (6,1,1) $67$
2 (8,0,1) $219$
2 (1,5,0) $10$
2 (3,4,0) $36$
2 (5,3,0) $132$
2 (7,2,0) $473$
2 (9,1,0) $1734$
2 (11,0,0) $6620$
Table 2: $g=0,M_{6}^{1},d=3$
d (a,b,c) $N^{0}_{d,a,b}$
3 (1,0,5) 0
3 (0,2,4) 11
3 (2,1,4) 55
3 (4,0,4) $30$
3 (1,3,3) 99
3 (3,2,3) $45$
3 (5,1,3) $225$
3 (7,0,3) $1011$
3 (0,5,2) $16$
3 (2,4,2) $76$
3 (4,3,2) $385$
3 (6,2,2) $1931$
3 (8,1,2) $9386$
3 (10,0,2) $45954$
3 (1,6,1) $128$
3 (3,5,1) $664$
3 (5,4,1) $3512$
3 (7,3,1) $18469$
3 (9,2,1) $96548$
3 (11,1,1) $511012$
3 (13,0,1) $2770596$
3 (0,8,0) $188$
3 (2,7,0) $1108$
3 (4,6,0) $6216$
3 (6,5,0) $34780$
3 (8,4,0) $194024$
3 (10,3,0) $1085892$
3 (12,2,0) $6165822$
3 (14,1,0) $35806494$
3 (16,0,0) $213709980$
Table 3: $g=0,M_{6}^{1},d=4$ part 1
d (a,b,c) $N^{0}_{d,a,b}$
4 (0,0,7) 11
4 (1,1,6) 99
4 (3,0,6) $61$
4 (0,3,5) $14$
4 (2,2,5) $107$
4 (4,1,5) $732$
4 (6,0,5) $4830$
4 (1,4,4) $178$
4 (3,3,4) $1218$
4 (5,2,4) $8133$
4 (7,1,4) $52507$
4 (9,0,4) $324764$
4 (0,6,3) $320$
4 (2,5,3) $2056$
4 (4,4,3) $13962$
4 (6,3,3) $94104$
4 (8,2,3) $622980$
4 (10,1,3) $4063860$
4 (12,0,3) $26578256$
4 (1,7,2) $3516$
4 (3,6,2) $23968$
4 (5,5,2) $166936$
4 (7,4,2) $1159218$
4 (9,3,2) $7990720$
4 (11,2,2) $54948346$
4 (13,1,2) $380720598$
4 (15,0,2) $2679044142$
Table 4: $g=0,M_{6}^{1},d=4$ part 2
d (a,b,c) $N^{0}_{d,a,b}$
4 (0,9,1) $5552$
4 (2,8,1) $40492$
4 (4,7,1) $291632$
4 (6,6,1) $2110864$
4 (8,5,1) $15251816$
4 (10,4,1) $110031632$
4 (12,3,1) $796460052$
4 (14,2,1) $5823161346$
4 (16,1,1) $43242657488$
4 (18,0,1) $327439797532$
4 (1,10,0) $63740$
4 (3,9,0) $493976$
4 (5,8,0) $3748804$
4 (7,7,0) $28346212$
4 (9,6,0) $213984472$
4 (11,5,0) $1617593360$
4 (13,4,0) $12302188692$
4 (15,3,0) $94605276228$
4 (17,2,0) $738764469204$
4 (19,1,0) $5876564125104$
4 (21,0,0) $47723447905060$
Table 5: $g=0,M_{6}^{1},d=5$ part 1
d (a,b,c) $N^{0}_{d,a,b}$
5 (0,1,8) $10$
5 (2,0,8) $161$
5 (1,2,7) $246$
5 (3,1,7) $2390$
5 (5,0,7) $20670$
5 (0,4,6) $432$
5 (2,3,6) $3915$
5 (4,2,6) $34180$
5 (6,1,6) $284685$
5 (8,0,6) $2269330$
5 (1,5,5) $6700$
5 (3,4,5) $57200$
5 (5,3,5) $484345$
5 (7,2,5) $4001415$
5 (9,1,5) $32175350$
5 (11,0,5) $252923350$
5 (0,7,4) $11980$
5 (2,6,4) $97660$
5 (4,5,4) $829884$
5 (6,4,4) $7042024$
5 (8,3,4) $58977314$
5 (10,2,4) $487020090$
5 (12,1,4) $3986631790$
5 (14,0,4) $32664263244$
5 (1,8,3) $168160$
5 (3,7,3) $1426788$
5 (5,6,3) $12341640$
5 (7,5,3) $106742892$
5 (9,4,3) $917273760$
5 (11,3,3) $7835510640$
5 (13,2,3) $66838183448$
5 (15,1,3) $572970400800$
5 (17,0,3) $4963870717184$
Table 6: $g=0,M_{6}^{1},d=5$ part 2
d (a,b,c) $N^{0}_{d,a,b}$
5 (0,10,2) $275340$
5 (2,9,2) $2427884$
5 (4,8,2) $21445040$
5 (6,7,2) $190810312$
5 (8,6,2) $1697371800$
5 (10,5,2) $15059634800$
5 (12,4,2) $133462672144$
5 (14,3,2) $1185922233290$
5 (16,2,2) $10614063989964$
5 (18,1,2) $96073499325220$
5 (20,0,2) $882272821107200$
5 ($1,1$) $3941780$
5 (3,10,1) $36486980$
5 (5,9,1) $335462284$
5 (7,8,1) $3085793380$
5 (9,7,1) $28364597480$
5 (11,6,1) $260604570680$
5 (13,5,1) $2398306990560$
5 (15,4,1) $22179879220568$
5 (17,3,1) $206859628175260$
5 (19,2,1) $1951736958419580$
5 (21,1,1) $18676573063528460$
5 (23,0,1) $181610832693333060$
5 (0,13,0) $5953000$
5 (2,12,0) $59294040$
5 (4,11,0) $573878820$
5 (6,10,0) $5489009900$
5 (8,9,0) $52316386080$
5 (10,8,0) $497999093480$
5 (12,7,0) $4743580571280$
5 (14,6,0) $45328575942720$
5 (16,5,0) $435813009759000$
5 (18,4,0) $4228201646521080$
5 (20,3,0) $41500751630424420$
5 (22,2,0) $412968880889100580$
5 (24,1,0) $4173087641902059600$
5 (26,0,0) $42876778851631702000$
Table 7: $g=1,M_{6}^{1},d=1,2$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
1 (0,1,1) $\frac{5}{24}$ 0
1 (2,0,1) $\frac{5}{24}$ 0
1 (1,2,0) $\frac{5}{12}$ 0
1 (3,1,0) $\frac{5}{8}$ 0
1 (5,0,0) $\frac{25}{24}$ 0
2 (1,0,3) $\frac{3}{8}$ 0
2 (0,2,2) $\frac{3}{8}$ 0
2 (2,1,2) $\frac{17}{12}$ 0
2 (4,0,2) $\frac{91}{24}$ 0
2 (1,3,1) $\frac{25}{12}$ 0
2 (3,2,1) $\frac{173}{24}$ 0
2 (5,1,1) $\frac{545}{24}$ 0
2 (7,0,1) $\frac{1765}{24}$ 0
2 (0,5,0) $\frac{10}{3}$ 0
2 (2,4,0) $\frac{73}{6}$ 0
2 (4,3,0) $\frac{89}{2}$ 0
2 (6,2,0) $\frac{3805}{24}$ 0
2 (8,1,0) $\frac{1735}{3}$ 0
2 (10,0,0) 2200-2200 0
Table 8: $g=1,M_{6}^{1},d=3$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
3 (0,0,5) 0 0
3 (1,1,4) $\frac{55}{24}$ 0
3 (3,0,4) $\frac{55}{4}$ 0
3 (0,3,3) $\frac{95}{24}$ 0
3 (2,2,3) $\frac{475}{24}$ 0
3 (4,1,3) $\frac{785}{8}$ 0
3 (6,0,3) $\frac{10337}{24}$ 11
3 (1,4,2) $\frac{65}{2}$ 0
3 (3,3,2) $\frac{3935}{24}$ 0
3 (5,2,2) $\frac{6519}{8}$ 11
3 (7,1,2) $\frac{15603}{4}$ $14$
3 (9,0,2) $\frac{75273}{4}$ $114$
3 (0,6,1) $\frac{160}{3}$ 0
3 (2,5,1) $\frac{830}{3}$ 0
3 (4,4,1) $\frac{4373}{3}$ 22
3 (6,3,1) $\frac{182575}{24}$ $25$
3 (8,2,1) $\frac{236081}{6}$ $222$
3 (10,1,1) $\frac{1236305}{6}$ $1650$
3 (12,0,1) $\frac{2212985}{2}$ $11325$
3 (1,7,0) $\frac{2705}{6}$ 0
3 (3,6,0) $\frac{7600}{3}$ 55
3 (5,5,0) $\frac{84815}{6}$ $55$
3 (7,4,0) $\frac{235232}{3}$ $468$
3 (9,3,0) $\frac{2613469}{6}$ $3558$
3 (11,2,0) $\frac{29454205}{12}$ $25275$
3 (13,1,0) $\frac{56619575}{4}$ $173490$
3 (15,0,0) $\frac{503847475}{6}$ $1175300$
Table 9: $g=1,M_{6}^{1},d=4$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
4 (0,1,6) $\frac{35}{8}$ 0
4 (2,0,6) $\frac{725}{24}$ 0
4 (1,2,5) $\frac{1255}{24}$ 0
4 (3,1,5) $\frac{2165}{6}$ 0
4 (5,0,5) $\frac{28625}{12}$ 0
4 (0,4,4) $\frac{341}{4}$ 11
4 (2,3,4) $\frac{6991}{12}$ 44
4 (4,2,4) $\frac{31035}{8}$ $32$
4 (6,1,4) $\frac{594955}{24}$ $310$
4 (8,0,4) $\frac{449893}{3}$ $3220$
4 (1,5,3) $\frac{2881}{3}$ $14$
4 (3,4,3) $\frac{25957}{4}$ $96$
4 (5,3,3) $\frac{173685}{4}$ $785$
4 (7,2,3) $\frac{1135543}{4}$ $6755$
4 (9,1,3) 1819093-1819093 $57960$
4 (11,0,3) $\frac{70041271}{6}$ $473586$
4 (0,7,2) $\frac{9647}{6}$ $29$
4 (2,6,2) 10889-10889 $228$
4 (4,5,2) 75377-75377 $1799$
4 (6,4,2) $\frac{6227887}{12}$ $14745$
4 (8,3,2) $\frac{10604846}{3}$ $122155$
4 (10,2,2) $\frac{287430247}{12}$ $1004916$
4 (12,1,2) $\frac{1961004899}{12}$ $8127687$
4 (14,0,2) $\frac{13594466647}{12}$ $65017656$
4 (1,8,1) $\frac{107875}{6}$ $460$
4 (3,7,1) $\frac{386698}{3}$ $3837$
4 (5,6,1) $\frac{2782210}{3}$ $31470$
4 (7,5,1) $\frac{19937800}{3}$ $259325$
4 (9,4,1) 47440464-47440464 $2139212$
4 (11,3,1) $\frac{1017676760}{3}$ $17589245$
4 (13,2,1) $\frac{9795944647}{4}$ $144007483$
4 (15,1,1) $\frac{53891135695}{3}$ $1178586170$
4 (17,0,1) $\frac{806789380135}{6}$ $9691893740$
4 (0,10,0) $\frac{164635}{6}$ $870$
4 (2,9,0) $\frac{639692}{3}$ $7708$
4 (4,8,0) $\frac{9683117}{6}$ $65124$
4 (6,7,0) $\frac{72844595}{6}$ $543660$
4 (8,6,0) $\frac{272974555}{3}$ $4535180$
4 (10,5,0) $\frac{2045100320}{3}$ $37854490$
4 (12,4,0) $\frac{10265421909}{2}$ $316190712$
4 (14,3,0) $\frac{78131149107}{2}$ $2646562486$
4 (16,2,0) $\frac{604014726315}{2}$ $22264309750$
4 (18,1,0) 2379763852305-2379763852305 $188886527100$
4 (20,0,0) $\frac{114958360544525}{6}$ $1620988570200$
Table 10: $g=1,M_{6}^{1},d=5$, part 1
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
5 (1,0,8) $\frac{1945}{24}$ 33
5 (0,2,7) $\frac{1435}{12}$ 88
5 (2,1,7) $\frac{2405}{2}$ $45$
5 (4,0,7) $\frac{126757}{12}$ $324$
5 (1,3,6) $\frac{15315}{8}$ $105$
5 (3,2,6) $\frac{50707}{3}$ $771$
5 (5,1,6) $\frac{3386155}{24}$ $6330$
5 (7,0,6) $\frac{2231555}{2}$ $55935$
5 (0,5,5) 3205-3205 $220$
5 (2,4,5) $\frac{54877}{2}$ $1694$
5 (4,3,5) $\frac{5573975}{24}$ $14090$
5 (6,2,5) $\frac{15246635}{8}$ $123855$
5 (8,1,5) $\frac{181156655}{12}$ $1127945$
5 (10,0,5) $\frac{1390208755}{12}$ $10460910$
5 (1,6,4) $\frac{274369}{6}$ $3401$
5 (3,5,4) $\frac{1160363}{3}$ $29084$
5 (5,4,4) $\frac{6522287}{2}$ $257413$
5 (7,3,4) $\frac{324143825}{12}$ $2335035$
5 (9,2,4) $\frac{2632411415}{12}$ $21477895$
5 (11,1,4) $\frac{7029536823}{4}$ $198302628$
5 (13,0,4) $\frac{84374784907}{6}$ $1821344223$
5 (0,8,3) 77296-77296 $6224$
5 (2,7,3) $\frac{1297759}{2}$ $56623$
5 (4,6,3) 5569635-5569635 $513600$
5 (6,5,3) $\frac{143175565}{3}$ $4719945$
5 (8,4,3) $\frac{1214792630}{3}$ $43829370$
5 (10,3,3) $\frac{6804949031}{2}$ $409413777$
5 (12,2,3) $\frac{85400501522}{3}$ $3832247481$
5 (14,1,3) $\frac{1434207179585}{6}$ $35880218130$
5 (16,0,3) 2028186128080-2028186128080 $336649015020$
5 (1,9,2) $\frac{2158487}{2}$ $104214$
5 (3,8,2) $\frac{28335680}{3}$ $986530$
5 (5,7,2) 83306285-83306285 $9296155$
5 (7,6,2) $\frac{2200182005}{3}$ $87964690$
5 (9,5,2) 6423114302-6423114302 $836194988$
5 (11,4,2) 56047183707-56047183707 $7977161338$
5 (13,3,2) $\frac{5873452441955}{12}$ $76321377460$
5 (15,2,2) 4301607795510-4301607795510 $732727776780$
5 (17,1,2) $\frac{229390883916695}{6}$ $7072862332690$
5 (19,0,2) $\frac{1034728934410532}{3}$ $68819539583286$
Table 11: $g=1,M_{6}^{1},d=5$, part 2
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
5 (0,11,1) $\frac{3402365}{2}$ $184480$
5 (2,10,1) $\frac{94016375}{6}$ $1825740$
5 (4,9,1) $\frac{857749505}{6}$ $17812920$
5 (6,8,1) $\frac{7821853175}{6}$ $172988900$
5 (8,7,1) 11859338793-11859338793 $1681583302$
5 (10,6,1) $\frac{322851405685}{3}$ $16388989480$
5 (12,5,1) $\frac{2929215576475}{3}$ $160253018950$
5 (14,4,1) 8891388713520-8891388713520 $1573356927950$
5 (16,3,1) $\frac{489643030178765}{6}$ $15531173703960$
5 (18,2,1) $\frac{4546452758023601}{6}$ $154438870292624$
5 (20,1,1) $\frac{14277641483246075}{2}$ $1550155293988500$
5 (22,0,1) $\frac{410343935038064375}{6}$ $15736008036054900$
5 (1,12,0) $\frac{73969625}{3}$ $3274890$
5 (3,11,0) $\frac{1428317155}{6}$ $33189090$
5 (5,10,0) $\frac{13579189585}{6}$ $332653440$
5 (7,9,0) $\frac{64202869786}{3}$ $3320681718$
5 (9,8,0) 201753545535-201753545535 $33152681950$
5 (11,7,0) 1900031299700-1900031299700 $331728574700$
5 (13,6,0) 17926222775700-17926222775700 $3331531685250$
5 (15,5,0) $\frac{510047851228460}{3}$ $33628906542030$
5 (17,4,0) $\frac{4879475383947517}{3}$ $341728487604966$
5 (19,3,0) $\frac{31485553789132145}{2}$ $3501911868774390$
5 (21,2,0) $\frac{927227199825166475}{6}$ $36252054632249100$
5 (23,1,0) $\frac{4624335742996347470}{3}$ $379702695213071370$
5 (25,0,0) 15644867027685574170-15644867027685574170 $4029226126511838330$
Table 12: $g=0,M_{6}^{2},d=1,2,3$
d (a,b,c) $N^{0}_{d,a,b}$
1 (0,1,1) 44
1 (2,0,1) 44
1 (1,2,0) 88
1 (3,1,0) $12$
1 (5,0,0) 2020
2 (0,0,3) 88
2 (1,1,2) $16$
2 (3,0,2) $32$
2 (0,3,1) $16$
2 (2,2,1) $64$
2 (4,1,1) $184$
2 (6,0,1) $576$
2 (1,4,0) $96$
2 (3,3,0) $368$
2 (5,2,0) $1280$
2 (7,1,0) $4632$
2 (9,0,0) $17704$
3 (1,0,4) $64$
3 (0,2,3) $64$
3 (2,1,3) $320$
3 (4,0,3) $1152$
3 (1,3,2) $448$
3 (3,2,2) $2240$
3 (5,1,2) $9888$
3 (7,0,2) $45504$
3 (0,5,1) $640$
3 (2,4,1) $3712$
3 (4,3,1) $19552$
3 (6,2,1) $98272$
3 (8,1,1) $505552$
3 (10,0,1) $2702000$
3 (1,6,0) $6144$
3 (3,5,0) $35584$
3 (5,4,0) $199424$
3 (7,3,0) $1102752$
3 (9,2,0) $6213728$
3 (11,1,0) $36112336$
3 (13,0,0) $217541136$
Table 13: $g=0,M_{6}^{2},d=4$
d (a,b,c) $N^{0}_{d,a,b}$
4 (0,1,5) $384$
4 (2,0,5) $2560$
4 (1,2,4) $3328$
4 (3,1,4) $20096$
4 (5,0,4) $103296$
4 (0,4,3) $4608$
4 (2,3,3) $31616$
4 (4,2,3) $198912$
4 (6,1,3) $1177152$
4 (8,0,3) $7139008$
4 (1,5,2) $48640$
4 (3,4,2) $345088$
4 (5,3,2) $2312960$
4 (7,2,2) $15125760$
4 (9,1,2) $100512384$
4 (11,0,2) $687547904$
4 (0,7,1) $74240$
4 (2,6,1) $579584$
4 (4,5,1) $4263424$
4 (6,4,1) $30403072$
4 (8,3,1) $215004288$
4 (10,2,1) $1541202944$
4 (12,1,1) $11310265600$
4 (14,0,1) $85345355904$
4 (1,8,0) $980992$
4 (3,7,0) $7649280$
4 (5,6,0) $58502144$
4 (7,5,0) $441611520$
4 (9,4,0) $3335082240$
4 (11,3,0) $25524363520$
4 (13,2,0) $199433493504$
4 (15,1,0) $1597307381568$
4 (17,0,0) $13143513966080$
Table 14: $g=0,M_{6}^{2},d=5$, part 1
d (a,b,c) $N^{0}_{d,a,b}$
5 (0,0,7) $3072$
5 (1,1,6) $33280$
5 (3,0,6) $238592$
5 (0,3,5) $45056$
5 (2,2,5) $357376$
5 (4,1,5) $2592000$
5 (6,0,5) $17329920$
5 (1,4,4) $530432$
5 (3,3,4) $4331520$
5 (5,2,4) $33241088$
5 (7,1,4) $246732544$
5 (9,0,4) $1859635968$
5 (0,6,3) $774144$
5 (2,5,3) $6976512$
5 (4,4,3) $59192320$
5 (6,3,3) $482989056$
5 (8,2,3) $3885555456$
5 (10,1,3) $31625447936$
5 (12,0,3) $263294153728$
5 (1,7,2) $11044864$
5 (3,6,2) $101627904$
5 (5,5,2) $898333696$
5 (7,4,2) $7760710656$
5 (9,3,2) $66707246848$
5 (11,2,2) $579410893568$
5 (13,1,2) $5127561935232$
5 (15,0,2) $46417851808512$
5 (0,9,1) $17539072$
5 (2,8,1) $171974656$
5 (4,7,1) $1620875264$
5 (6,6,1) $14922719232$
5 (8,5,1) $135757207552$
5 (10,4,1) $1235286731776$
5 (12,3,1) $11356463308416$
5 (14,2,1) $106142531433856$
5 (16,1,1) $1012183022176064$
5 (18,0,1) $9868286056471104$
Table 15: $g=0,M_{6}^{2},d=5$, part 2
d (a,b,c) $N^{0}_{d,a,b}$
5 (1,10,0) $295157760$
5 (3,9,0) $2892357632$
5 (5,8,0) $28004556800$
5 (7,7,0) $268191540224$
5 (9,6,0) $2557415162880$
5 (11,5,0) $24477957189120$
5 (13,4,0) $236783110502912$
5 (15,3,0) $2325894666007680$
5 (17,2,0) $23271693753656448$
5 (19,1,0) $237640853649062080$
5 (21,0,0) $2479850169230342720$
Table 16: $g=1,M_{6}^{2},d=1,2,3$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
1 (1,0,1) $\frac{1}{2}$ 0
1 (0,2,0) 1-1 0
1 (2,1,0) $\frac{3}{2}$ 0
1 (4,0,0) $\frac{5}{2}$ 0
2 (0,1,2) $\frac{10}{3}$ 0
2 (2,0,2) $\frac{20}{3}$ 0
2 (1,2,1) $\frac{40}{3}$ 0
2 (3,1,1) $\frac{115}{3}$ 0
2 (5,0,1) 120-120 0
2 (0,4,0) 20-20 0
2 (2,3,0) $\frac{230}{3}$ 0
2 (4,2,0) $\frac{800}{3}$ 0
2 (6,1,0) 965-965 0
2 (8,0,0) $\frac{11065}{3}$ 0
3 (0,0,4) $\frac{56}{3}$ 0
3 (1,1,3) 88-88 0
3 (3,0,3) $\frac{928}{3}$ 0
3 (0,3,2) 120-120 0
3 (2,2,2) 600-600 0
3 (4,1,2) $\frac{7852}{3}$ 0
3 (6,0,2) 11936-11936 0
3 (1,4,1) $\frac{2944}{3}$ 0
3 (3,3,1) $\frac{15476}{3}$ 0
3 (5,2,1) 25756-25756 0
3 (7,1,1) $\frac{395158}{3}$ 88
3 (9,0,1) $\frac{2101942}{3}$ $136$
3 (0,6,0) $\frac{4832}{3}$ 0
3 (2,5,0) $\frac{27952}{3}$ 0
3 (4,4,0) 52112-52112 0
3 (6,3,0) $\frac{861020}{3}$ $16$
3 (8,2,0) 1611044-1611044 $272$
3 (10,1,0) 9333366-9333366 $3000$
3 (12,0,0) $\frac{168264818}{3}$ $27240$
Table 17: $g=1,M_{6}^{2},d=4$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
4 (1,0,5) $\frac{2560}{3}$ 0
4 (0,2,4) 1056-1056 0
4 (2,1,4) 6320-6320 0
4 (4,0,4) $\frac{94192}{3}$ $256$
4 (1,3,3) 9680-9680 0
4 (3,2,3) $\frac{181024}{3}$ $256$
4 (5,1,3) $\frac{1050152}{3}$ $3584$
4 (7,0,3) 2088664-2088664 $30720$
4 (0,5,2) $\frac{43520}{3}$ 0
4 (2,4,2) $\frac{308864}{3}$ $512$
4 (4,3,2) $\frac{2053856}{3}$ $6144$
4 (6,2,2) $\frac{13260032}{3}$ $59392$
4 (8,1,2) $\frac{87089248}{3}$ $499296$
4 (10,0,2) $\frac{589841888}{3}$ $4007808$
4 (1,6,1) 169856-169856 $1280$
4 (3,5,1) $\frac{3742112}{3}$ $12800$
4 (5,4,1) 8838784-8838784 $118784$
4 (7,3,1) $\frac{185860784}{3}$ $1043936$
4 (9,2,1) 440296032-440296032 $8781120$
4 (11,1,1) 3206552104-3206552104 $72704112$
4 (13,0,1) $\frac{72109405664}{3}$ $602772128$
4 (0,8,0) 283776-283776 $3584$
4 (2,7,0) 2210240-2210240 $28672$
4 (4,6,0) 16860672-16860672 $251904$
4 (6,5,0) $\frac{379892480}{3}$ $2211200$
4 (8,4,0) $\frac{2850607936}{3}$ $19140736$
4 (10,3,0) 7225953840-7225953840 $163776288$
4 (12,2,0) $\frac{168399937952}{3}$ $1399658624$
4 (14,1,0) $\frac{1341957809680}{3}$ $12047200912$
4 (16,0,0) 3665151541000-3665151541000 $105004257360$
Table 18: $g=1,M_{6}^{2},d=5$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
5 (0,1,6) 11840-11840 0
5 (2,0,6) 84224-84224 0
5 (1,2,5) 120832-120832 $768$
5 (3,1,5) $\frac{2583392}{3}$ $13824$
5 (5,0,5) $\frac{16612064}{3}$ $192512$
5 (0,4,4) $\frac{518144}{3}$ $2048$
5 (2,3,4) $\frac{4207744}{3}$ $28416$
5 (4,2,4) $\frac{31740928}{3}$ $328960$
5 (6,1,4) $\frac{229497440}{3}$ $3457152$
5 (8,0,4) $\frac{1689720416}{3}$ $32359808$
5 (1,5,3) 2195328-2195328 $60928$
5 (3,4,3) 18503296-18503296 $649728$
5 (5,3,3) $\frac{446421376}{3}$ $6655104$
5 (7,2,3) 1174715232-1174715232 $65456768$
5 (9,1,3) 9391653184-9391653184 $618276096$
5 (11,0,3) 76961465920-76961465920 $5753289984$
5 (0,7,2) 3375616-3375616 $132096$
5 (2,6,2) 31075840-31075840 $1330176$
5 (4,5,2) 272878080-272878080 $13367296$
5 (6,4,2) $\frac{6989383808}{3}$ $132810240$
5 (8,3,2) $\frac{59226474208}{3}$ $1299074944$
5 (10,2,2) 169095332896-169095332896 $12544392576$
5 (12,1,2) $\frac{4432072954096}{3}$ $120900975936$
5 (14,0,2) $\frac{39663151820800}{3}$ $1173261629056$
5 (1,8,1) $\frac{154276864}{3}$ $2738176$
5 (3,7,1) $\frac{1451733632}{3}$ $27183616$
5 (5,6,1) $\frac{13282234496}{3}$ $271901184$
5 (7,5,1) $\frac{119682815168}{3}$ $2715012608$
5 (9,4,1) $\frac{1076978890048}{3}$ $27008532736$
5 (11,3,1) $\frac{9791608439120}{3}$ $268391724096$
5 (13,2,1) $\frac{90566832590384}{3}$ $2679621302336$
5 (15,1,1) $\frac{855469605426472}{3}$ $27017537730048$
5 (17,0,1) $\frac{8269323668574136}{3}$ $276138960259456$
5 (0,10,0) 86736896-86736896 $5566464$
5 (2,9,0) $\frac{2547003392}{3}$ $55035904$
5 (4,8,0) $\frac{24584536064}{3}$ $555792384$
5 (6,7,0) 78025121792-78025121792 $5650489344$
5 (8,6,0) $\frac{2214589877504}{3}$ $57577844736$
5 (10,5,0) 7003208209024-7003208209024 $587949802496$
5 (12,4,0) 67142653348736-67142653348736 $6030797709312$
5 (14,3,0) $\frac{1961908485684176}{3}$ $62362747054848$
5 (16,2,0) 6492457837633744-6492457837633744 $652295164595968$
5 (18,1,0) 65834214232016696-65834214232016696 $6919499749525632$
5 (20,0,0) 682743285735788648-682743285735788648 $74585887608233472$
Table 19: $g=0,M_{6}^{3},d=1,2,3$
d (a,b,c) $N^{0}_{d,a,b}$
1 (1,0,1) $18$
1 (0,2,0) $45$
1 (2,1,0) $63$
1 (4,0,0) $108$
2 (1,0,2) $108$
2 (0,2,1) $378$
2 (2,1,1) $864$
2 (4,0,1) $2754$
2 (1,3,0) $2187$
2 (3,2,0) $7047$
2 (5,1,0) $25758$
2 (7,0,0) $102060$
3 (1,0,3) $1944$
3 (0,2,2) $7452$
3 (2,1,2) $24300$
3 (4,0,2) $110808$
3 (1,3,1) $65610$
3 (3,2,1) $290142$
3 (5,1,1) $1469664$
3 (7,0,1) $8019000$
3 (0,5,0) $131220$
3 (2,4,0) $715149$
3 (4,3,0) $3766014$
3 (6,2,0) $21210984$
3 (8,1,0) $126574812$
3 (10,0,0) $796767840$
Table 20: $g=0,M_{6}^{3},d=4$
d (a,b,c) $N^{0}_{d,a,b}$
4 (1,0,4) $62208$
4 (0,2,3) $248832$
4 (2,1,3) $1057536$
4 (4,0,3) $6216912$
4 (1,3,2) $2991816$
4 (3,2,2) $16726176$
4 (5,1,2) $107372952$
4 (7,0,2) $741917880$
4 (0,5,1) $6495390$
4 (2,4,1) $42554646$
4 (4,3,1) $278396352$
4 (6,2,1) $1955545416$
4 (8,1,1) $14516081280$
4 (10,0,1) $112889878200$
4 (1,6,0) $94242204$
4 (3,5,0) $683570907$
4 (5,4,0) $4991884362$
4 (7,3,0) $38081408688$
4 (9,2,0) $303386439924$
4 (11,1,0) $2518696831248$
4 (13,0,0) $21763790077104$
Table 21: $g=0,M_{6}^{3},d=5$
d (a,b,c) $N^{0}_{d,a,b}$
5 (1,0,5) $2916000$
5 (0,2,4) $11955600$
5 (2,1,4) $62694000$
5 (4,0,4) $449017344$
5 (1,3,3) $183760488$
5 (3,2,3) $1238349384$
5 (5,1,3) $9579584880$
5 (7,0,3) $79760239920$
5 (0,5,2) $423735624$
5 (2,4,2) $3238483356$
5 (4,3,2) $25212033432$
5 (6,2,2) $211216698360$
5 (8,1,2) $1868651427024$
5 (10,0,2) $17269190883936$
5 (1,6,1) $7459305876$
5 (3,5,1) $63063741510$
5 (5,4,1) $543094438248$
5 (7,3,1) $4895543320344$
5 (9,2,1) $46024322567256$
5 (11,1,1) $449367253012512$
5 (13,0,1) $4544618405916240$
5 (0,8,0) $15633399897$
5 (2,7,0) $145384720047$
5 (4,6,0) $1335091394610$
5 (6,5,0) $12453847443000$
5 (8,4,0) $119983325683644$
5 (10,3,0) $1195990165100736$
5 (12,2,0) $12325456056749400$
5 (14,1,0) $131218747829213400$
5 (16,0,0) $1442401273691663040$
Table 22: $g=1,M_{6}^{3},d=1,2,3$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
1 (0,0,1) $\frac{9}{4}$ 0
1 (1,1,0) $\frac{63}{8}$ 0
1 (3,0,0) $\frac{27}{2}$ 0
2 (0,0,2) 18-18 0
2 (1,1,1) $\frac{297}{2}$ 0
2 (3,0,1) $\frac{1917}{4}$ 0
2 (0,3,0) $\frac{2943}{8}$ 0
2 (2,2,0) $\frac{9657}{8}$ 0
2 (4,1,0) $\frac{8883}{2}$ 0
2 (6,0,0) $\frac{35235}{2}$ 0
3 (0,0,3) 378-378 $27$
3 (1,1,2) $\frac{9963}{2}$ $81$
3 (3,0,2) 22842-22842 $243$
3 (0,3,1) $\frac{54351}{4}$ $81$
3 (2,2,1) $\frac{239841}{4}$ $486$
3 (4,1,1) 303993-303993 $2187$
3 (6,0,1) 1659690-1659690 $10935$
3 (1,4,0) $\frac{1186083}{8}$ $729$
3 (3,3,0) $\frac{3120849}{4}$ $4374$
3 (5,2,0) 4394898-4394898 $24057$
3 (7,1,0) $\frac{52463943}{2}$ $137781$
3 (9,0,0) 165166614-165166614 $826686$
Table 23: $g=1,M_{6}^{3},d=4$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
4 (0,0,4) 12393-12393 $2187$
4 (1,1,3) 232065-232065 $13851$
4 (3,0,3) 1366875-1366875 $67797$
4 (0,3,2) 684774-684774 $18954$
4 (2,2,2) 3743496-3743496 $143613$
4 (4,1,2) 23908041-23908041 $918540$
4 (6,0,2) 164904903-164904903 $6259194$
4 (1,4,1) $\frac{38738331}{4}$ $253692$
4 (3,3,1) $\frac{125225433}{2}$ $1972674$
4 (5,2,1) 437451759-437451759 $14541363$
4 (7,1,1) 3237427035-3237427035 $110257605$
4 (9,0,1) 25123505859-25123505859 $868020300$
4 (0,6,0) $\frac{42972363}{2}$ $470205$
4 (2,5,0) $\frac{1239605451}{8}$ $3965760$
4 (4,4,0) $\frac{4495376313}{4}$ $32264811$
4 (6,3,0) $\frac{17075109663}{2}$ $259323525$
4 (8,2,0) $\frac{135676682451}{2}$ $2124189360$
4 (10,1,0) 562258758393-562258758393 $17881926768$
4 (12,0,0) 4854224373084-4854224373084 $155016025290$
Table 24: $g=1,M_{6}^{3},d=5$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
5 (0,0,5) 538488-538488 $195372$
5 (1,1,4) 13584186-13584186 $2038284$
5 (3,0,4) 97309836-97309836 $13288212$
5 (0,3,3) 43113789-43113789 $3446712$
5 (2,2,3) 278158212-278158212 $30018033$
5 (4,1,3) 2127258450-2127258450 $237683160$
5 (6,0,3) 17629551342-17629551342 $1984125132$
5 (1,4,2) $\frac{1505585475}{2}$ $59337684$
5 (3,3,2) 5719059675-5719059675 $538946784$
5 (5,2,2) 47352243843-47352243843 $4765188690$
5 (7,1,2) 416074383666-416074383666 $43330628592$
5 (9,0,2) 3826973405736-3826973405736 $407674059804$
5 (0,6,1) 1749516894-1749516894 $118262025$
5 (2,5,1) $\frac{58224030795}{4}$ $1152035055$
5 (4,4,1) 123411045312-123411045312 $11006956674$
5 (6,3,1) 1101441692466-1101441692466 $104895765531$
5 (8,2,1) 10285610571117-10285610571117 $1020477653388$
5 (10,1,1) 99935930525508-99935930525508 $10198147299480$
5 (12,0,1) 1006928375152050-1006928375152050 $104869613909880$
5 (1,7,0) $\frac{269759628747}{8}$ $2441617101$
5 (3,6,0) 306334343151-306334343151 $24628038822$
5 (5,5,0) 2827530502056-2827530502056 $247068849834$
5 (7,4,0) $\frac{54059816716407}{2}$ $2496139936101$
5 (9,3,0) 267911657815050-267911657815050 $25688164768974$
5 (11,2,0) 2749583878851825-2749583878851825 $270571180960692$
5 (13,1,0) 29185732286251029-29185732286251029 $2922525674615178$
5 (15,0,0) 320196437236815048-320196437236815048 $32395182807076992$
Table 25: $g=0,M_{6}^{4}$
d (a,b,c) $N^{0}_{d,a,b}$
1 (0,0,1) $96$
1 (1,1,0) $416$
1 (3,0,0) $736$
2 (0,1,1) $5568$
2 (2,0,1) $21120$
2 (1,2,0) $65536$
2 (3,1,0) $249856$
2 (5,0,0) $1050368$
3 (1,0,2) $534528$
3 (0,2,1) $1572864$
3 (2,1,1) $9142272$
3 (4,0,1) $53127168$
3 (1,3,0) $29884416$
3 (3,2,0) $173998080$
3 (5,1,0) $1098065920$
3 (7,0,0) $7453751296$
4 (0,0,3) $13160448$
4 (1,1,2) $294912000$
4 (3,0,2) $2247081984$
4 (0,3,1) $918552576$
4 (2,2,1) $7183269888$
4 (4,1,1) $56891523072$
4 (6,0,1) $471610097664$
4 (1,4,0) $24075304960$
4 (3,3,0) $188432777216$
4 (5,2,0) $1580113821696$
4 (7,1,0) $14045898735616$
4 (9,0,0) $132127907905536$
5 (0,1,3) $8889827328$
5 (2,0,3) $84506443776$
5 (1,2,2) $263901413376$
5 (3,1,2) $2562163605504$
5 (5,0,2) $25399993171968$
5 (0,4,1) $848390258688$
5 (2,3,1) $8334908325888$
5 (4,2,1) $83350340763648$
5 (6,1,1) $867295299895296$
5 (8,0,1) $9412512794935296$
5 (1,5,0) $28347455242240$
5 (3,4,0) $278636624084992$
5 (5,3,0) $2909478586155008$
5 (7,2,0) $31928942775238656$
5 (9,1,0) $367130295597662208$
5 (11,0,0) $4420079855887319040$
Table 26: $g=1,M_{6}^{4}$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
1 (0,1,0) $\frac{260}{3}$ 0
1 (2,0,0) $\frac{460}{3}$ 0
2 (1,0,1) 4864-4864 0
2 (0,2,0) $\frac{43744}{3}$ 0
2 (2,1,0) $\frac{167008}{3}$ 0
2 (4,0,0) $\frac{698720}{3}$ 0
3 (0,0,2) 132864-132864 $768$
3 (1,1,1) 2246144-2246144 $13440$
3 (3,0,1) 13061376-13061376 $73344$
3 (0,3,0) $\frac{21873152}{3}$ $27136$
3 (2,2,0) 42461696-42461696 $161280$
3 (4,1,0) $\frac{804326144}{3}$ $960512$
3 (6,0,0) $\frac{5455746304}{3}$ $5898240$
4 (0,1,2) 74569728-74569728 $2571264$
4 (2,0,2) 571607040-571607040 $18837504$
4 (1,2,1) 1818378240-1818378240 $55492608$
4 (3,1,1) 14434416640-14434416640 $422731776$
4 (5,0,1) 119608152064-119608152064 $3456921600$
4 (0,4,0) $\frac{18345644032}{3}$ $146182144$
4 (2,3,0) $\frac{143574781952}{3}$ $1153077248$
4 (4,2,0) $\frac{1204880285696}{3}$ $9562345472$
4 (6,1,0) $\frac{10713421205504}{3}$ $83229306880$
4 (8,0,0) 33606505377792-33606505377792 $754787155968$
5 (1,0,3) 21040939008-21040939008 $2005917696$
5 (0,2,2) 65329790976-65329790976 $6142722048$
5 (2,1,2) 638763810816-638763810816 $57498402816$
5 (4,0,2) 6343662845952-6343662845952 $561661526016$
5 (1,3,1) 2081253425152-2081253425152 $175519629312$
5 (3,2,1) 20872275427328-20872275427328 $1716412514304$
5 (5,1,1) 217273901776896-217273901776896 $17729712807936$
5 (7,0,1) 2356423499333632-2356423499333632 $191826913124352$
5 (0,5,0) $\frac{21453605306368}{3}$ $517300158464$
5 (2,4,0) $\frac{210823503020032}{3}$ $5103810117632$
5 (4,3,0) $\frac{2201894952304640}{3}$ $53278992793600$
5 (6,2,0) $\frac{24166727101579264}{3}$ $582424512397312$
5 (8,1,0) $\frac{277963197651501056}{3}$ $6634027667816448$
5 (10,0,0) 1116531188802797568-1116531188802797568 $78493526034874368$
Table 27: $g=0,M_{6}^{5}$
d (a,b,c) $N^{0}_{d,a,b}$
1 (0,1,0) $3250$
1 (2,0,0) $6125$
2 (0,0,1) $247500$
2 (1,1,0) $3718750$
2 (3,0,0) $17406875$
3 (1,0,1) $659250000$
3 (0,2,0) $2700512500$
3 (2,1,0) $19190225000$
3 (4,0,0) $150549428125$
4 (0,1,1) $529823250000$
4 (2,0,1) $4729124250000$
4 (1,2,0) $20567866625000$
4 (3,1,0) $209913851312500$
4 (5,0,0) $2337181124531250$
5 (0,0,2) $109236016800000$
5 (1,1,1) $5092187634000000$
5 (3,0,1) $61252356251250000$
5 (0,3,0) $21811124012125000$
5 (2,2,0) $274577525136875000$
5 (4,1,0) $3688919538904687500$
5 (6,0,0) $53412041211701171875$
Table 28: $g=1,M_{6}^{5}$
d (a,b,c) $N^{1}_{d,a,b}$ $E_{d,a,b,c}$
1 (1,0,0) $\frac{18375}{8}$ 0
2 (0,1,0) $\frac{8038625}{6}$ 0
2 (2,0,0) $\frac{148120375}{24}$ 0
3 (0,0,1) 246708750-246708750 $947500$
3 (1,1,0) 7059310625-7059310625 $14139375$
3 (3,0,0) $\frac{1319375639375}{24}$ $93667500$
4 (1,0,1) 1762031831250-1762031831250 $34818450000$
4 (0,2,0) $\frac{23015099134375}{3}$ $111635590625$
4 (2,1,0) $\frac{155974421459375}{2}$ $1070162215625$
4 (4,0,0) $\frac{5194583816496875}{6}$ $10767522628125$
5 (0,1,1) 1870863418500000-1870863418500000 $97094656425000$
5 (2,0,1) 22353636563531250-22353636563531250 $1193292296250000$
5 (1,2,0) $\frac{303154762779953125}{3}$ $4376885918562500$
5 (3,1,0) $\frac{8130487814000265625}{6}$ $56589153313015625$
5 (5,0,0) $\frac{470853639798192671875}{24}$ $767616850277828125$
Appendix A Enumeration of Degree $3$ Elliptic Curves in $M_{6}^{k}$ ($k=1,2,3,4,5$)
It is well-known that any degree $3$ elliptic curve in $CP^{N-1}$ is given as a **plane cubic**. Therefore, we first consider $Gr(3,N)$, the moduli space of $2$-planes in $CP^{N-1}$. A cubic curve in a plane (biholomorphic to $CP^{2}$) is determined by its defining equation, a degree $3$ homogeneous polynomial in three homogeneous coordinates of $CP^{2}$. Thus, the moduli space of degree $3$ elliptic curves in $CP^{N-1}$ is given by the fiber space $\pi_{P}:P(S^{3}U^{*})\rightarrow Gr(3,N)$, where $\pi_{U}:U\rightarrow Gr(3,N)$ is the tautological rank $3$ vector bundle over $Gr(3,N)$.
Next, we construct the moduli space of elliptic curves in $M_{6}^{k}$ for $k$ varying from $1$ to $5$.
  • **Case k=1k=1:** Since $M_{6}^{1}$ is biholomorphic to $CP^{4}$, the moduli space is $\pi_{P}:P(S^{3}U^{*})\rightarrow Gr(3,5)$.
  • **Case k=2k=2:** For a degree $3$ elliptic curve in $CP^{5}$ to be contained in $M_{6}^{2}$, the $2$-plane that contains the elliptic curve must itself be contained in $M_{6}^{2}$. Hence, the moduli space is given by $\pi_{P}^{-1}(PD(c_{\text{top}}(S^{2}U^{*})))$, where $PD(\alpha)$ is the submanifold of $Gr(3,6)$ that is Poincaré dual to the cohomology class $\alpha\in H^{*}(Gr(3,6))$, and $\pi_{P}:P(S^{3}U^{*})\rightarrow Gr(3,6)$ is the projection of the fiber space.
  • **Cases k=3,4,5k=3,4,5:** For a degree $3$ elliptic curve in $CP^{5}$ to be contained in $M_{6}^{k}$, the restriction of the defining equation of $M_{6}^{k}$ (a homogeneous degree $k$ polynomial in $6$ homogeneous coordinates of $CP^{5}$) to the $2$-plane containing the elliptic curve must be divisible by the defining equation of the elliptic curve. Let $\pi_{S}:S_{P}\rightarrow P(S^{3}U^{*})$ be the tautological line bundle of the projectivization $P(S^{3}U^{*})$. Using this line bundle and the aforementioned condition, the moduli space of degree $3$ elliptic curves in $M_{6}^{k}$ is given by $PD(c_{\text{top}}(S^{k}U^{*}/(S^{k-3}U^{*}\otimes S_{P})))$.
Next, we briefly discuss how to express the condition that the elliptic curve passes through $PD(h^{a})\subset CP^{N-1}$ in terms of $H^{*}(P(S^{3}U^{*}))$. For this purpose, we introduce the moduli space $\mathcal{M}$ of degree $3$ elliptic curves with one marked point. Let $\pi_{F}:\mathcal{M}\rightarrow P(S^{3}U^{*})$ be the forgetful map that omits the marked point, and let $ev:\mathcal{M}\rightarrow CP^{N-1}$ be the evaluation map. With these setups, the passing-through condition is represented by $\pi_{F*}(ev^{*}(h^{a}))\in H^{a-1}(P(S^{3}U^{*}))$, where $\pi_{F*}:H^{m}(\mathcal{M})\rightarrow H^{m-1}(P(S^{3}U^{*}))$ denotes fiber integration (or push-forward) by $\pi_{F}$. For details on the construction of $\mathcal{M}$, we recommend readers refer to [11]. The cohomology class $\pi_{F*}(ev^{*}(h^{a}))$ is determined using the well-known projection formula [3], but we omit the computational details as they are highly technical. Instead, we present the explicit form of $\pi_{F*}(ev^{*}(h^{a}))$ at the end of this appendix.
Finally, we introduce the generators and relations of $H^{*}(P(S^{3}U^{*}))$ for explicit computation. Let $c_{i}$ ($i=1,2,3$) be the $i$-th Chern class of the vector bundle $U^{*}$. $H^{*}(Gr(3,N))$ is generated by these classes, and the relations are given as follows. Let $h_{i}(c_{1},c_{2},c_{3})$ be the weighted homogeneous polynomial of degree $i$ defined by
$\frac{1}{1-c_{1}t+c_{2}t^{2}-c_{3}t^{3}}=1+\sum_{i=1}^{\infty}h_{i}t^{i}.$ (A.11)
Then, we have
$H^{*}(Gr(3,N))\simeq{\bf C}[c_{1},c_{2},c_{3}]/(h_{N-2},h_{N-1},h_{N}),$ (A.12)
where $(h_{N-2},h_{N-1},h_{N})$ is the ideal of ${\bf C}[c_{1},c_{2},c_{3}]$ generated by $h_{N-2},h_{N-1}$, and $h_{N}$. We now describe the ring structure of $H^{*}(P(S^{3}U^{*}))$. For this purpose, we introduce another generator $z$, which plays the role of the hyperplane class of the fiber projective space. Since $S^{3}U^{*}$ is a rank $10$ vector bundle over $Gr(3,N)$, $z$ satisfies the relation
$R(z,c_{1},c_{2},c_{3}):=z^{10}+\sum_{j=1}^{10}c_{i}(S^{3}U^{*})z^{10-i}=0.$ (A.13)
Hence, the ring structure of $H^{*}(P(S^{3}U^{*}))$ is given by
$H^{*}(P(S^{3}U^{*}))\simeq{\bf C}[c_{1},c_{2},c_{3},z]/(h_{N-2},h_{N-1},h_{N},R).$ (A.14)
The integration rule for cohomology elements of $H^{*}(P(S^{3}U^{*}))$ is fixed by the following equality:
$\int_{P(S^{3}U^{*})}z^{9}\cdot(c_{3})^{N-3}=1.$ (A.15)
The explicit forms of $\pi_{F*}(ev^{*}(h^{a}))$ for $a=2,3,4$ are given as follows:
$\pi_{F*}(ev^{*}(h^{2}))$ $=$ $3c_{1}+z=:G_{1}$
$\pi_{F*}(ev^{*}(h^{3}))$ $=$ $3c_{1}^{2}-3c_{2}+c_{1}z=:G_{2}$
$\pi_{F*}(ev^{*}(h^{4}))$ $=$ $3c_{1}^{3}-6c_{1}c_{2}+3c_{3}+(c_{1}^{2}-c_{2})z=:G_{3}.$
These expressions do not depend on $N$. Let $\tilde{E}_{3,a,b,c}$ be the number of degree $3$ elliptic curves in $M_{6}^{k}$ that corresponds to $E_{3,a,b,c}$ in our conjecture. With these preparations, $\tilde{E}_{3,a,b,c}$ is explicitly computed by the following formulae:
$\tilde{E}_{3,a,b,c}$ $=$ $\int_{P(S^{3}U^{*})}G_{1}^{a}G_{2}^{b}G_{3}^{c}\quad(k=1,\;N=5),$
$\tilde{E}_{3,a,b,c}$ $=$ $\int_{P(S^{3}U^{*})}c_{\text{top}}(S^{2}U^{*})G_{1}^{a}G_{2}^{b}G_{3}^{c}\quad(k=2,\;N=6),$
$\tilde{E}_{3,a,b,c}$ $=$ $\int_{P(S^{3}U^{*})}c_{\text{top}}(S^{k}U^{*}/(S^{k-3}U^{*}\otimes S_{P}))G_{1}^{a}G_{2}^{b}G_{3}^{c}\quad(k=3,4,5,\;N=6).$
We find that $\tilde{E}_{3,a,b,c}$ indeed coincides with $E_{3,a,b,c}$  [^2] A PDF copy of the Maple worksheet used for computing $\tilde{E}_{3,a,b,c}$ of $M_{6}^{5}$ is available on M. Jinzenji’s ResearchGate homepage [9].
Acknowledgment
We would like to thank Prof. H. Suzuki for valuable discussions.
Funding
Our research is partially supported by JSPS Grant No. 22K03289. The research of M. J. is also partially supported by JSPS Grant No. 24H00182.
Disclosure Statements on Competing Interests
We have no relevant financial or non-financial ompeting interests to declare.